plz show all work...

Exercises for Chapter 7: Sa mpling Distributions I entrance exam at an MBA program in the Central Valley of doo Are scores on the GMAT entrance exam a discrete or a continuous variable? What is the probability that a randomly selected application will report a GMAT score of less than 600? a. b.

estith that a sample of 50 randomly selected applications will report an average GMAT score of less than 600? d. W hat is the probability GMAT score of less than 600? IAT score of les t100 randomly selected applications will report an average e warehouse for a hardw g the shipment, the purchasing manager insists that a sample of the drills be randomly selected are store chain has just received a truckload of 5000 electric drills. for testing. He intends to measure if the mean f e pu the maximum power consumption of each drill. He will reject the shipment umption for the sample is greater than 300 watts as listed on the product label. he drills on the truck Accompanying the shipment is the manufacturers records of product tests showing t require an average of 295 watts with a standard deviation of consumption for population of drills. 16 watts. Assume normal distribution of power a. Find the probability that the truckload of drill will be rejected if he samples 25 drills. b. Find the probability that the truckload of drills will be rejected if he samples 49 drills c. What are the impacts of larger sample size n on the mean and standard deviation among sample averages of power consumption (Ax and ơx) and its absolute values of z-score 121 ? Exercise 7.3: The life span of Good Old Everglo Bulbs follows a normal distribution with a mean life of 400 hours and a standard deviation of 30 hours. Assume infinite population of Good Old Everglo Bulbs. What is the probability that 4 bulbs selected at random will have an average life span of more than 445 hours?
Answer #1

Solution:-

7.1)

Mean = 591, S.D = 42

a) The scores on GMAT entrance exam are continuous variable.

b) The probability that scores will be less than 600 is 0.5848.

x = 600

By applying normal distribution:-

z = \frac{x-\mu }{\sigma }

z = 0.2143

P(z < 0.2143) = 0.5848

c) The probability that scores from sample of 50 will be less than 600 is 0.9351.

x = 600

By applying normal distribution:-

z = \frac{x-\mu }{\frac{\sigma }{\sqrt{n}}}

z = 1.515

P(z < 1.515) = 0.9351

d) The probability that scores from sample of 100 will be less than 600 is 0.9839.

x = 600

By applying normal distribution:-

z = \frac{x-\mu }{\frac{\sigma }{\sqrt{n}}}

z = 2.143

P(z < 2.143) = 0.9839